# Strong-to-weak spontaneous breaking of 1-form symmetry and intrinsically mixed topological order

### v2

@@ Appendix A Two-way channel between completely mixed state and SW-SSB
[…]
@@ Appendix B Choi state TEE and LW CMI
$\begin{split}&\gamma(\rho)\\ &=-\log\frac{\sum_{\sigma_{n}\in C\cup D,\sigma_{m}\in A\cup D}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\rho|\sigma_{n,u}\sigma_{n,l}^{\dagger}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\rho|\sigma_{m,u}\sigma_{m,l}^{\dagger}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}}{\sum_{\sigma_{n}\in A\cup C\cup D,\sigma_{m}\in D}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\rho|\sigma_{n,u}\sigma_{n,l}^{\dagger}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\rho|\sigma_{m,u}\sigma_{m,l}^{\dagger}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}}\end{split}$ (116)
Like in (LABEL:I2choi), we find that $\gamma(\rho)$ can be written in terms of expectation values of operators in the Choi state. Let us first evaluate $\gamma(\rho)$ when $|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}=|tc\rangle|tc\rangle$ (i.e. $\rho=|tc\rangle\langle tc|$), where $|tc\rangle$ is the fixed point toric code wavefunction. In this case, the only nonzero expectation values are those of closed $X_{u}X_{l}$ and $Z_{u}Z_{e}$ loops. These operators have expectation value 1. The number of closed $X_{u}X_{l}$ loops in a simply connected region $R$ is $2^{N_{v,R}}$ while the number of $Z_{u}Z_{e}$ loops is $2^{N_{p,R}}$. It follows that
@@
We can also compute $\gamma(\rho)$ easily in the SW-SSB phase. For the fixed point state $|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\rho|=4\prod_{e}\left(\frac{1+Z_{e,u}Z_{e,l}}{2}\right)\prod_{v}\left(\frac{1+A_{v,u}A_{v,l}}{2}\right)\prod_{p}\left(\frac{1+B_{p,u}}{2}\right)$, the only operators with nonzero expectation value are products (not necessarily closed loops) of $Z_{e,u}Z_{e,l}$ and closed $X_{e,u}X_{e,l}$ loops. We find that
@@ Appendix C Rényi-2 local indistinguishable sectors and SW-SSB
On the torus, we can define $W_{x},W_{y},T_{x},$ and $T_{y}$, which are the $\mathbb{Z}_{2}$ Wilson and ’t-Hooft 1-form symmetry operators along non-contractible cycles in the $x/y$ directions. It is well known that in an isolated quantum system, the toric code Hamiltonian has four locally indistinguishable ground states on a torus, coming from complete spontaneous breaking of this $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ 1-form symmetry. In particular, at the toric code fixed point, all four non-contractible loop operators commute with the Hamiltonian. We can choose a basis for the ground state space that consists of eigenstates of $W_{x}$ and $W_{y}$, so $W_{x}|\psi_{1}\rangle=\alpha_{x}|\psi_{1}\rangle$, $W_{y}|\psi_{1}\rangle=\alpha_{y}|\psi_{1}\rangle$. However, due to the commutation relations between $W_{y}$ and $T_{x}$, there is another state $T_{y}|\psi_{1}\rangle=|\psi_{2}\rangle$ that has eigenvalue $-1$ under $W_{x}$: $W_{x}(T_{y}|\psi_{1}\rangle)=-\alpha T_{y}|\psi_{1}\rangle=-\alpha|\psi_{2}\rangle$. Similarly, we can apply $T_{y}$ and $T_{x}T_{y}$ on $|\psi_{1}\rangle$ to get two other states orthogonal to $|\psi_{1}\rangle$. In total, we get four orthogonal states on a torus, labeled by their different eigenvalues under $W_{x}$ and $W_{y}$: $\{|\psi_{i}\rangle\}=\{|\psi_{1}\rangle,T_{x}|\psi_{1}\rangle,T_{y}|\psi_{1}\rangle,T_{x}T_{y}|\psi_{1}\rangle\}$. These states form a basis for the ground state space of the toric code Hamiltonian. Importantly, the above four states are locally indistinguishable:
@@
for any local operator $O$ [^20] Clearly $\langle\psi_{i}|T_{x}|\psi_{j}\rangle\neq c_{T_{x}}\delta_{ij}$ because $T_{x}$ is a global operator. Here, $c_{O}$ is a constant: the expectation value of any local operator is the same in each of the ground states. The local indistinguishability of the toric code ground states is what allows them to form a robust quantum code.
@@ Appendix D Decohered toric code
[…]
@@
$|\rho_{0}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}=|\psi_{0}\rangle|\psi_{0}^{*}\rangle\propto\sum_{g_{u},g_{l}}|g_{u},g_{l}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}},$⟩ , (127)
@@
$|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}$⟩ $=e^{\mu\sum Z_{u}Z_{e}}|\rho_{0}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}\propto\sum_{g_{u},g_{l}}e^{-2\mu|g_{u}+g_{l}|}|g_{u},g_{l}\rangle,$ (128)
@@
$|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}\propto|\psi_{0}\rangle\sum_{h}e^{-2\mu|h|}|h\rangle.$ (129)
@@
$\begin{split}&\frac{\mathrm{Tr}(\rho\prod_{e\in{\gamma}}Z_{e}\rho\prod_{e^{\prime}\in{\gamma}}Z_{e^{\prime}})}{\mathrm{Tr}\rho^{2}}=\frac{\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\rho|\prod_{e\in{\gamma}}Z_{e}^{u}Z_{e}^{l}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}}{\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\rho|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}}\\ =&\frac{\sum_{h}\langle h|\prod_{e\in\gamma}Z_{e}e^{-4\mu|h|}|h\rangle}{\sum_{h}\langle h|e^{-4\mu|h|}|h\rangle}\\ \end{split}.$ (130)
@@
$\begin{split}&\frac{\mathrm{Tr}(\rho\prod_{e\in\tilde{\gamma}}X_{e}\mathcal{E}[\rho_{0}]\prod_{e^{\prime}\in\tilde{\gamma}}X_{e^{\prime}})}{\mathrm{Tr}\rho^{2}}\\ =&\frac{\mathrm{Tr}(\rho\mathcal{E}[\prod_{e\in\tilde{\gamma}}X_{e}\rho_{0}\prod_{e^{\prime}\in\tilde{\gamma}}X_{e^{\prime}}])}{\mathrm{Tr}\rho^{2}}=\frac{\mathrm{Tr}(\rho\mathcal{E}[\rho_{0}])}{\mathrm{Tr}\rho^{2}}=1.\\ \end{split}$ (131)
@@
where $O_{i(j)}$ is an operator fully supported on regions $i$ ($j$), $v_{LR}$ and $\mu$ are constants that depend on locality and norms of local terms in the Lindbladian $\mathcal{L}$ and $|i-j|$ is the distance between $i$ and $j$. Here $||O||$ is the operator norm defined as $||O||=\sup_{\|\psi\|=1}||O\psi||.$ This means that local Lindbladian evolution can never cause a diverging correlation length for linear correlation functions, since $\mathrm{Tr}(\rho O)\leq\|O\|$ for an apropriately normalized $\rho$. Therefore, we do not expect linear order and disorder parameters to change in scaling.
@@ D.1 Rényi-2 Markov length
×$\times\left(\sum_{\sigma^{\prime}\in\bar{C}}\prod_{v}A_{v}^{\sigma_{v}^{\prime}}(1-2p_{v})^{N_{\sigma^{\prime}}}\right),$ (135)
@@
Appendix E Sequential circuit for $|\rho(\beta)\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}$⟩
@@
[…]
Figure 4: Unitary circuit relating $|\rho_{B}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}$⟩ to $|\rho_{\beta}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}$⟩ when we take one of the directions (say, $y$) to have open boundary conditions. (a) We can write $e^{\beta A_{v}}|\uparrow_{m}\rangle$ as $e^{i\tilde{\beta}\tilde{A}_{v,m}}|\uparrow_{m}\rangle$ where $\tilde{A}_{v,m}$ has a single $Y_{e,m}$ operator. (b) We can transform all the $A_{v,m}$ terms in a single row ($x$ direction) simultaneously. (c) Side view of the circuit. The circuit is a sequential circuit in the $y$ direction because the gates in different $y$ coordinates do not commute. The $Y_{e,m}$ operator of the $\tilde{A}_{v,m}$ one row below overlaps with a $X_{e,m}$ operator of the $\tilde{A}_{v,m}$ one row above. The naive construction illustrated here allows periodic boundary conditions in the $x$ direction but not in the $y$ direction; closing up the circuit in the $y$ direction would require some modifications.
@@
[…]
@@
Because the circuit is sequential in the $y$ direction, it gives a correlation length in the $y$ direction. In other words, $U(\beta)O_{e}U^{\dagger}(\beta)$ generically maps a local operator $O_{e}$ supported on site $e$ to an operator with strict locality in the $x$ direction but an exponential tail in the $+y$ direction (note that there is no tail in the $-y$ direction).
[…]

### v3

@@ Appendix A Two-way channel between completely mixed state and SW-SSB
We will show in this appendix that there is an explicit two-way finite depth quantum channel between the maximally mixed state and the SW-SSB state $\rho_{B}=\frac{2^{N_{p}}}{2^{N_{e}}}\prod_{p}\left(\frac{1+B_{p}}{2}\right)$ on a sphere. Getting from $\rho_{B}$ to the maximally mixed state is straightforward – one can simply apply the incoherent $X$ error channel, which is finite depth. To get from maximally mixed state to $\rho_{B}$, we first apply a complete amplitude damping channel to map to the product state $|\uparrow\rangle\langle\uparrow|$. It is instructive to first construct the complete amplitude damping channel for a single qubit. We want to map
@@ Appendix B Choi state TEE and LW CMI
$\begin{split}&\gamma(\rho)\\ &=-\log\frac{\sum_{\sigma_{n}\in C\cup D,\sigma_{m}\in A\cup D}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\rho|\sigma_{n,u}\sigma_{n,l}^{\dagger}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\rho|\sigma_{m,u}\sigma_{m,l}^{\dagger}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}}{\sum_{\sigma_{n}\in A\cup C\cup D,\sigma_{m}\in D}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\rho|\sigma_{n,u}\sigma_{n,l}^{\dagger}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\rho|\sigma_{m,u}\sigma_{m,l}^{\dagger}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}}\end{split}$ (116)
Like in (LABEL:I2choi), we find that $\gamma(\rho)$ can be written in terms of expectation values of operators in the Choi state. Let us first evaluate $\gamma(\rho)$ when $|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}=|tc\rangle|tc\rangle$ (i.e. $\rho=|tc\rangle\langle tc|$), where $|tc\rangle$ is the fixed point toric code wavefunction. In this case, the only nonzero expectation values are those of closed $X_{u}X_{l}$ and $Z_{u}Z_{e}$ loops. These operators have expectation value 1. The number of closed $X_{u}X_{l}$ loops in a simply connected region $R$ is $2^{N_{v,R}}$ while the number of $Z_{u}Z_{e}$ loops is $2^{N_{p,R}}$. It follows that
@@
We can also compute $\gamma(\rho)$ easily in the SW-SSB phase. For the fixed point state $|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\rho|=4\prod_{e}\left(\frac{1+Z_{e,u}Z_{e,l}}{2}\right)\prod_{v}\left(\frac{1+A_{v,u}A_{v,l}}{2}\right)\prod_{p}\left(\frac{1+B_{p,u}}{2}\right)$, the only operators with nonzero expectation value are products (not necessarily closed loops) of $Z_{e,u}Z_{e,l}$ and closed $X_{e,u}X_{e,l}$ loops. We find that
@@ Appendix C Rényi-2 local indistinguishable sectors and SW-SSB
On the torus, we can define $W_{x},W_{y},T_{x},$ and $T_{y}$, which are the $\mathbb{Z}_{2}$ Wilson and ’t-Hooft 1-form symmetry operators along non-contractible cycles in the $x/y$ directions. It is well known that in an isolated quantum system, the toric code Hamiltonian has four locally indistinguishable ground states on a torus, coming from complete spontaneous breaking of this $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ 1-form symmetry. In particular, at the toric code fixed point, all four non-contractible loop operators commute with the Hamiltonian. We can choose a basis for the ground state space that consists of eigenstates of $W_{x}$ and $W_{y}$, so $W_{x}|\psi_{1}\rangle=\alpha_{x}|\psi_{1}\rangle$, $W_{y}|\psi_{1}\rangle=\alpha_{y}|\psi_{1}\rangle$. However, due to the commutation relations between $W_{y}$ and $T_{x}$, there is another state $T_{y}|\psi_{1}\rangle=|\psi_{2}\rangle$ that has eigenvalue 1-1 under $W_{x}$: $W_{x}(T_{y}|\psi_{1}\rangle)=-\alpha T_{y}|\psi_{1}\rangle=-\alpha|\psi_{2}\rangle$. Similarly, we can apply $T_{y}$ and $T_{x}T_{y}$ on $|\psi_{1}\rangle$ to get two other states orthogonal to $|\psi_{1}\rangle$. In total, we get four orthogonal states on a torus, labeled by their different eigenvalues under $W_{x}$ and $W_{y}$: $\{|\psi_{i}\rangle\}=\{|\psi_{1}\rangle,T_{x}|\psi_{1}\rangle,T_{y}|\psi_{1}\rangle,T_{x}T_{y}|\psi_{1}\rangle\}$. These states form a basis for the ground state space of the toric code Hamiltonian. Importantly, the above four states are locally indistinguishable:
@@
for any local operator OO. [^20] Clearly $\langle\psi_{i}|T_{x}|\psi_{j}\rangle\neq c_{T_{x}}\delta_{ij}$ because $T_{x}$ is a global operator. Here, $c_{O}$ is a constant: the expectation value of any local operator is the same in each of the ground states. The local indistinguishability of the toric code ground states is what allows them to form a robust quantum code.
@@ Appendix D Decohered toric code
where $|\uparrow\rangle$ is the state with all spins in the +1+1 eigenstate of $Z$. $|\psi_{0}\rangle$ can be written as an equal superposition of all closed loop states $|g\rangle$, where $|g\rangle$ is a product state with spins in the 1-1 eigenstate of $Z_{e}$ only along the closed loops $g$:
@@
$|\rho_{0}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}=|\psi_{0}\rangle|\psi_{0}^{*}\rangle\propto\sum_{g_{u},g_{l}}|g_{u},g_{l}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}},$ (127)
@@
[…]
@@
$|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}\propto|\psi_{0}\rangle\sum_{h}e^{-2\mu|h|}|h\rangle.$ (129)
@@
$\begin{split}&\frac{\mathrm{Tr}(\rho\prod_{e\in{\gamma}}Z_{e}\rho\prod_{e^{\prime}\in{\gamma}}Z_{e^{\prime}})}{\mathrm{Tr}\rho^{2}}=\frac{\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\rho|\prod_{e\in{\gamma}}Z_{e}^{u}Z_{e}^{l}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}}{\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\rho|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}}\\ =&\frac{\sum_{h}\langle h|\prod_{e\in\gamma}Z_{e}e^{-4\mu|h|}|h\rangle}{\sum_{h}\langle h|e^{-4\mu|h|}|h\rangle}\\ \end{split}.$ (130)
@@
[…]
@@
[…]
@@ D.1 Rényi-2 Markov length
$\times\left(\sum_{\sigma^{\prime}\in\bar{C}}\prod_{v}A_{v}^{\sigma_{v}^{\prime}}(1-2p_{v})^{N_{\sigma^{\prime}}}\right),$ (135)
@@
Appendix E Sequential circuit for $|\rho(\beta)\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$
@@
on a system with open boundary conditions. The trick we use is similar to that used in Ref. Jones et al., 2021: we pull $Z_{e}$ operators out from $|\uparrow\rangle$ to turn $\prod e^{\beta_{A}A_{v}}$ into a sequential circuit on this particular state. This sequential circuit is exact for open boundary conditions in at least one direction. Closing up the boundary seems a bit tricky though, at least if we want to stay exact. We will show that this circuit maps local operators to exponentially localized operators. We can use this unitary operator to obtain a quasilocal parent Hamiltonian that has a constant gap for all (finite) $\beta$ because it is just $H_{B}$ conjugated by a unitary operator.
Figure 4: Unitary circuit relating $|\rho_{B}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$ to $|\rho_{\beta}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$ when we take one of the directions (say, $y$) to have open boundary conditions. (a) We can write $e^{\beta A_{v}}|\uparrow_{m}\rangle$ as $e^{i\tilde{\beta}\tilde{A}_{v,m}}|\uparrow_{m}\rangle$ where $\tilde{A}_{v,m}$ has a single $Y_{e,m}$ operator. (b) We can transform all the $A_{v,m}$ terms in a single row ($x$ direction) simultaneously. (c) Side view of the circuit. The circuit is a sequential circuit in the $y$ direction because the gates in different $y$ coordinates do not commute. The $Y_{e,m}$ operator of the $\tilde{A}_{v,m}$ one row below overlaps with a $X_{e,m}$ operator of the $\tilde{A}_{v,m}$ one row above. The naive construction illustrated here allows periodic boundary conditions in the $x$ direction but not in the $y$ direction; closing up the circuit in the $y$ direction would require some modifications.
@@
Therefore, by pulling out $Z_{e}$ operators from $|\uparrow\rangle$, we can turn a single non-unitary gate $e^{\beta A_{v}}$ into a unitary operator $e^{i\tilde{\beta}\tilde{A}_{v}}$, which has the same action on the state $|\uparrow\rangle$ up to a proportionality constant. Now we need to perform this transformation on every site. Note that we can do this transformation in one direction (say, the $x$ direction) all at once (see Fig. 4.(b)), but then we need to do this transformation in the other direction (say, the $y$ direction) sequentially (see Fig. 4.(c)), because the gates after this transformation no longer commute. Therefore, we end up with a circuit that is sequential in one direction. We can have periodic boundary conditions in the $x$ direction, but modifying the circuit to have periodic boundary conditions in the $y$ direction seems trickier (at least for an exact circuit solution).
@@
Because the circuit is sequential in the $y$ direction, it gives a correlation length in the $y$ direction. In other words, $U(\beta)O_{e}U^{\dagger}(\beta)$ generically maps a local operator $O_{e}$ supported on site $e$ to an operator with strict locality in the $x$ direction but an exponential tail in the +y+y direction (note that there is no tail in the y-y direction).
Next, we must show that $U(\beta)Z_{e}U^{\dagger}(\beta)=\tilde{Z}_{e}$ is exponentially localized. Since $\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\rho_{B}|Z_{e}Z_{e^{\prime}}|\rho_{B}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}=1$, this would also show that $\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\rho_{\beta}|\tilde{Z}_{e}\tilde{Z}_{e^{\prime}}|\rho_{\beta}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}=1$, i.e. there is an exponentially localized dressed order parameter with expecation value 1.